OpenStudy (anonymous):

Anybody willing to grade my homework. I am willing to give free fan and medal.

OpenStudy (anonymous):

@Blacksteel

OpenStudy (anonymous):

@KamiBug may you two help grade it

OpenStudy (anonymous):

for me please

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

where's your homework ??

OpenStudy (anonymous):

Hold on trying to get it

OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 4 points) (2.06) Choose the correct description of the graph of the compound inequality: x - 3 < -9 or x + 5 greater than or equal to 12 A number line with a closed circle on -6, shading to the left, and an open circle on 7, shading to the right A number line with an open circle on -6, a closed circle on 7, and shading in between A number line with an open circle on -6, shading to the left, and a closed circle on 7, shading to the right A number line with a closed circle on -6, an open circle on 7, and shading in between Question 2 (Multiple Choice Worth 3 points) (2.03) Solve for x: 2|x - 3| + 1 = 7 x = 0, x = 6 x = -1, x = 6 x = 0, x = 9 No solutions Question 3 (Multiple Choice Worth 4 points) (2.06) The number line below represents which combined inequality? number line with a closed circle on -4, shading to the left and a closed circle on 3, shading to the right x greater than or equal to -4 and x less than or greater to 3 x greater than or equal to -4 or x less than or greater to 3 x less than or greater to -4 and x greater than or equal to 3 x less than or greater to -4 or x greater than or equal to 3 Question 4 (Multiple Choice Worth 4 points) (2.02) Sean, a freelance editor, charges the rates shown in the table below to edit manuscripts. The cost per page increases as the quality of editing improves. Sean also gives a 5% discount if the entire amount is paid up front. A 2-column table with 'Type of Editing' and 'Cost per page' as the column headings. Express Proofreading costs $2.00, Basic Proofreading - $2.95, Extended Proofreading - $3.95, and Deep Editing - $11.00 Michelle has an 85-page manuscript. The equation below shows the relationship between the total cost of editing 85 pages, T, and the cost per page, c, if she gets the 5% discount: 85c - 0.05(85c) = T Using the equation, what is the best quality of editing that Michelle can get done for a maximum of $161.50? Express proofreading Basic proofreading Extended proofreading Deep editing Question 5 (Multiple Choice Worth 3 points) (2.01) Solve: 6x - 2(x + 1) = 3x - 10 8 -8 -12 12 Question 6 (Multiple Choice Worth 3 points) (2.01) Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 5x - 2x = 2x + 2 5x + 2 = 2x - 10 The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describe solutions to equations shown. Question 7 (Multiple Choice Worth 4 points) (2.05) Solve the inequality 4x - 5 < 2x + 11. x > 8 x > -8 x < 8 x < -8 Question 8 (Multiple Choice Worth 4 points) (2.02) Choose the equation that could be used to find two consecutive integers whose sum is 69. n + (n + 2) = 69 n + (n + 1) = 69 n - (n + 1) = 69 n + 2n = 69 Question 9 (Multiple Choice Worth 3 points) (2.05) Choose the correct description of the graph of the inequality x - 3 > 5. Open circle on 8, shading to the left. Closed circle on 8, shading to the left. Open circle on 8, shading to the right. Closed circle on 8, shading to the right. Question 10 (Multiple Choice Worth 4 points) (2.07) Which of the following is not equivalent to the formula d = rt? d over r= t r = dt d over t= r d over r t= 1 Question 11 (Multiple Choice Worth 4 points) (2.06) Henry's rocket is designed to launch to a height of 100 feet in the air. The measurements so far have the launch height (h) varying no more than 4 feet. Which inequality below represents this scenario? |h - 4| less than or greater to 100 |h + 4| greater than or equal to 100 |h - 100| less than or greater to 4 |h + 100| greater than or equal to 4 Question 12 (Multiple Choice Worth 3 points) (2.01) After a discount of 10%, the price of a pair of jeans, including tax, was $69.62. The following equation can be used to find x, the price of the pair of jeans without the discount: x - 0.1x + 0.07(x - 0.1x) = 69.62 What is the price of the pair of jeans without the discount? $70.58 $72.29 $74.86 $80.02 Question 13 (Multiple Choice Worth 4 points) (2.06) Choose the correct description of the graph of the compound inequality: 2x + 5 > 7 and 5x less than or greater to 25 A number line with an open circle on 1, shading to the left, and a closed circle on 5, shading to the right A number line with an open circle on 1, a closed circle on 5, and shading in between A number line with a closed circle on 1, shading to the left, and an open circle on 5, shading to the right A number line with a closed circle on 1, an open circle on 5, and shading in between Question 14 (Multiple Choice Worth 3 points) (2.03) Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 |2x - 4| = 5 |3x - 4| = -12 Equation #1 and Equation #2 have the same number of solutions. Equation #1 has more solutions than equation #2. Equation #1 has fewer solutions than equation #2. None of the statements above describe the number of solutions to equations shown.

Nnesha (nnesha):

O_O

OpenStudy (texaschic101):

how are we supposed to grade it if we do not know what answers you picked

OpenStudy (anonymous):

1.c 2.a 3.d 4.d 5.b 6.b 7c 8.b 9.b 10.b 11,c 12.b 13.d 14.b

OpenStudy (anonymous):

@texaschic101 @KamiBug I have them right here

OpenStudy (anonymous):

Are you guys still here?

OpenStudy (texaschic101):

hold on I am checking them

OpenStudy (anonymous):

ok srry

OpenStudy (texaschic101):

9 is wrong

OpenStudy (texaschic101):

10 is wrong

OpenStudy (texaschic101):

I do not know # 11

OpenStudy (anonymous):

why are they wrong

OpenStudy (anonymous):

@KamiBug what do you think?

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

is number 9 d? and 10 c

OpenStudy (texaschic101):

I got B for # 11

OpenStudy (anonymous):

what do you guys think number 9 and 10 are

OpenStudy (texaschic101):

9 is C I was wrong about 10...it is B I am still checking the rest

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

I got B for 13

OpenStudy (texaschic101):

I am not sure about 14....but I agree with the majority of the answers you gave

OpenStudy (anonymous):

why did you get b for 11 and 13

OpenStudy (texaschic101):

I did not get B for 11.....I was probably looking at the wrong one when I said that...I do not know # 11.

OpenStudy (texaschic101):

#13. 2x + 5 > 7 2x > 7 - 5 2x > 2 x > 1 -- open circle (because there is no equal sign) shaded to right 5x <= 25 x < = 5 -- closed circle on 5, shaded to left so it is answer : B